OpenStudy (anonymous):

collect like terms 15m+7n-8m-9n

OpenStudy (anonymous):

7m-2n

OpenStudy (anonymous):

i have another questions for u

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

Find the slope , it exists of the line containing the pair of points (3,10) and (4,-4)

OpenStudy (anonymous):

you take (y1-y2)/(x1-x2) so (10-(-4))/(3-4)=(14/-1)=-14

OpenStudy (anonymous):

thanks i have 1 more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve using the multiplication principles . -1/6x=-9/10

jimthompson5910 (jim_thompson5910):

multiply both sides by 6 and reduce: x = 54/10 x = 27/5

jimthompson5910 (jim_thompson5910):

*multiply both sides by -6*

OpenStudy (anonymous):

have time for some more questions

jimthompson5910 (jim_thompson5910):

sure, just post to left

OpenStudy (anonymous):

ok